OpenStudy (anonymous):

Simplify the expression Please. 6 + (14 – 6) • 3 + 2 • 3= i Will Give Medal

OpenStudy (opcode):

Order of operation: BODMAS

OpenStudy (opcode):

@Jackieblue61 please don't just give the answer.

OpenStudy (anonymous):

no its pemdas right?

OpenStudy (anonymous):

parenthesis, exponents, multiply, divide, add, subtract. 14- 6 = 8... 8 x 3 = 24+2+6 = 36

OpenStudy (anonymous):

medal bro?

OpenStudy (e.mccormick):

P vs B = parenthesis vs braces, not much of a defference.

OpenStudy (opcode):

^ I grew up and learned that '(' and ')' were called soft/round brackets.

OpenStudy (e.mccormick):

Yes, that is British English. ( ) — round brackets, open brackets, brackets (UK), or parentheses (US) [ ] — square brackets, closed brackets, or brackets (US) { } — curly brackets, definite brackets, swirly brackets, birdie brackets, Scottish brackets, squirrelly brackets, braces, gullwings, fancy brackets, squiggly brackets, or curly braces (US)